OpenStudy (anonymous):

solve following inequality and write in interval notation: l3x-9l is > or = to 6

OpenStudy (anonymous):

\[\left| 3x-9 \right| \ge 6\]

OpenStudy (shadowlegendx):

\[\left| 3x - 9 \right|\ge 6\] Add 9 to both sides \[\left| 3x \right|\ge 15\] Divide both sides by 3 \[\left| x \right|\ge 5\]

OpenStudy (shadowlegendx):

@mando06 do you get it? :P

OpenStudy (anonymous):

what about interval notation

OpenStudy (shadowlegendx):

Omg I'm stupid, this is absolute value haha

OpenStudy (shadowlegendx):

So this type of absolute value inequality is called a disjunction. \[\left| x \right|> y\] \[\left| x \right| \ge y\] These are disjunctions

OpenStudy (shadowlegendx):

Disjunctions = or The way you set up this problem is simple. |3x−9|≥6 Convert to \[|3x−9|≥6 or |3x−9|\le -6\]

OpenStudy (shadowlegendx):

"or" is not a variable nor a set of variables, it's merely a divider that states that this equation is a disjunction.

OpenStudy (shadowlegendx):

*inequality

OpenStudy (shadowlegendx):

So now just solve both inequalities similar to how I did it before

OpenStudy (shadowlegendx):

|3x−9|≥6 Add 9 to both sides |3x|≥15 Divide both sides by 3 |x|≥5 |3x−9|≤−6 Add 9 to both sides |3x|≤ 3 Divide both sides by 3 |x|≤ 1

OpenStudy (shadowlegendx):

@mando06

OpenStudy (shadowlegendx):

\[\left| x \right| \ge 5 or \left| x \right| \le 1\]